OpenStudy (anonymous):

L varies inversely of P+5. If L=1 when P=4, find L when P=1.

OpenStudy (callisto):

L varies inversely of P+5, L(P+5) = constant 1(4+5) = L(1+5) 1(9) =6L L = 3/2

OpenStudy (anonymous):

Thanks for the help!